# ZipCache: A DRAM/SSD Cache with Built-in Transparent Compression

### v1

@@ Abstract
1 Introduction
@@
Key-value cache plays a crucial role in providing high-throughput, low-latency data services. Major Internet service providers, such as Google and Meta, often deploy a fleet of cache servers as the first line of defense to handle a massive influx of requests for key-value data, a type of unstructured data organized in simple forms as keys and values (e.g., “User ID” and “User name”). Key-value caching can accelerate data retrievals and alleviate the traffic to backend databases by serving from high-speed storage medium, typically DRAM, flash memory, or a combination of both, such as Meta’s CacheLib [2] and Ximalaya’s xcache [11].
@@
The effectiveness of key-value caching hinges on its ability of accommodating the requested data in cache. While a larger cache intuitively leads to improved performance, expanding the capacity of key-value caches within data centers is often not simply a matter of hardware upgrade. Many practical constraints, such as server costs, cooling expenses, rack space, real estate limitations, and even human resource expenses must be taken into consideration. Let us consider the hardware cost as an example: Microsoft Azure reports that DRAM constitutes 50% of their server costs [47], and Meta reports a similar trend (40% of the rack cost) [45]; Although the prices of high-speed NVMe SSDs are comparatively lower, they are still rather substantial [58, 48, 14]. Relying solely on hardware investment to increase cache capacity is apparently not a sustainable, cost-effective approach to keep up with the rapid growth of data. This poses an increasingly severe challenge in today’s data centers.
@@
A potential solution is compression. By condensing data to occupy a smaller footprint, one could virtually expand the cache capacity, allowing cache to hold more data, which in turn increases the cache hit ratio. Despite the adoption of data compression in computing systems in prior studies [24, 30, 51, 62], interestingly, this simple idea has not gained much traction in key-value cache design. We believe that this lack of adoption in practice is due to several unique and critical issues inherent in key-value cache systems:
@@
$\bullet$ Issue #1: The commonly used hash indexing causes random, compression-unfriendly data placement. Most key-value cache systems adopt a hash index based structure to manage the key-value data [6, 43, 20]. With hash indexing, keys are randomly dispersed in a flat, shallow data structure, which is advantageous for quick search in a large key space but comes with a detrimental effect for compression: Due to the nature of hash functions, the keys are evenly distributed, leaving unrelated data randomly mingled together. Such data layout is inherently difficult for effective data compression, as compression algorithms heavily rely on organizing similar data content within a close proximity.
@@
$\bullet$ Issue #2: The structure designed for managing small-size key-value items introduces a severe read/write amplification problem. Key-value workloads are known to be dominated by small-size data items. According to a study from Meta/Facebook, the majority of key-value items are (much) smaller than 500 bytes [15]. Since compressing each individual small key-value item yields limited or no benefits in size reduction, achieving effective compression requires to pack a collection of small key-value items for a reasonable compression ratio. However, this “optimization” would result in a substantial, undesirable increase in access operations, i.e., read/write amplification, when reading or updating a small key-value item in a much larger compression unit.
@@
We have implemented a prototype of ZipCache, which is a hybrid key-value cache with two cache layers, a DRAM layer and a flash memory layer. We use ScaleFlux’s CSD3000 SSD [8] with hardware acceleration for transparent on-device compression. Our experiments show promising results. In comparison to the state-of-the-art solutions, including CacheLib [18] and xcache [11] and Kangaroo [46], our evaluation results demonstrate that, with a design carefully optimized for data compression, ZipCache can achieve up to 72.4% higher throughput and 42.4% lower 90-percentile read latency, and reduces the SSD write amplification by up to 26.2$\times$. It is our hope that this work will motivate more future research to explore the full potential of the long-overlooked block compression for performance-critical caching systems.
@@
2 Background and Motivation
2.1 Data Compression
General-purpose block compression is realized by deduplicating repeated byte strings in a data block, referred to as LZ search [64, 65]. Although CPU-based LZ search suffers from low speed due to the high CPU cache miss rate, its reverse process (i.e., LZ decompression) can be much faster. The compression block size affects the trade-off between compression ratio [^1] In this work, we define compression ratio as $\frac{S_{orig}}{S_{comp}}\geq 1$, where $S_{orig}$ and $S_{comp}$ denote the size of the original and compressed data blocks. and (de)compression speed. Using the file samba in Silesia corpus [9] as test data, Fig. 1 shows the LZ4 compression ratio and (de)compression latency under different block sizes. It shows about 4$\times$ speed performance difference between compression and decompression. As the compression block size continues to increase, the compression ratio first significantly improves and then gradually saturates.
Figure 1: Comparison of compression ratio, (de)compression latency under different compression block size.
The decompression process scans through the LZ-compressed byte stream to sequentially reconstruct the original data block. In theory, this process could terminate at any byte location, leading to a partially reconstructed data block. This makes it possible to realize early termination of decompression: Suppose LZ search compresses an $n$-segment data block ${\bf D}=[b_{1},b_{2},\cdots,b_{n}]$ into ${\bf C}$. When decompressing ${\bf C}$, the original data $[b_{1},\cdots,b_{k}]$ are successively reconstructed as $k$ grows from 1 to $n$. Let $\tau$ denote the latency of decompressing ${\bf C}$ to reconstruct the entire ${\bf D}$. If the decompression process terminates once after the first $m$ segments $[b_{1},\cdots,b_{m}]$, where $m\leq n$) have been reconstructed, we define its early termination factor $\gamma(m,n)=m/n$, and the decompression latency is about $\gamma(m,n)\cdot\tau$. Suppose we are only interested in obtaining one cache object in the $m$-th segment, we could reduce the latency by $1-m/n$ via decompression early termination. Using file samba in Silesia corpus [9] as test data, we partitioned each 4KB data block into 256B ($n=16$) segments and measured the average LZ4 decompression latency under different early termination factor $\gamma(m,n)$ as shown in Table 1, which reveals substantial performance benefits.
Table 1: Decompression latency under different $\gamma(m,n)$.
@@ 2.1 Data Compression
2.2 Cache Index Data Structure
Most in-memory data stores (e.g., Redis [6], FASTER [20], MICA [43]) use hash index to reduce the latency and simplify the implementation. In contrast, most storage-based data stores (e.g., RocksDB [54, 28], WiredTiger [10], Bw-tree [40]) employ tree index to reduce the index memory usage and embrace the storage block I/O interface. As for hybrid-DRAM/SSD caches, Meta’s CacheLib [18, 2] uses hash index for the DRAM and SSD tiers, while Redis-compatible xcache [11] (developed based on Redis [6]) and Pika/RocksDB [5] employs hash index for DRAM tier and log-structured merge (LSM) tree [50] index for SSD tier.
The index structure has a substantial effect on the efficacy of data compression. Since compression ratio is proportional to the byte content similarity, a tree index that sorts all the cache objects based on their keys is clearly more beneficial, in comparison to hash index that randomly hashes cache objects into data blocks. For the purpose of demonstration, Fig. 2(a) shows the 4KB block compression ratio under hash index and B+ tree index. The key is 8B unix timestamp and object size ranging from 4B to 64B, which are both extracted from the Bitstamp Exchange Data [1]. By keeping objects with adjacent keys together, B+ tree achieves over 2$\times$ higher compression ratio than hash index. Although B+ tree index has a longer traversing latency than hash index, decompression tends to dominate the overall data access latency, which largely reduces the overall latency gap between B+ tree and hash index as shown in Fig. 2(b). A strong implication is that, although prior work on in-memory cache design widely adopted hash index, integrating block compression into in-memory cache makes tree index a favorable choice.
Figure 2: (a) Compression ratio of 4KB blocks under hash index and B+ tree index, and (b) latency of index traversing and block decompression under different total number of cache objects (hence different B+ tree depth), where key and values are obtained from Bitstamp Exchange Data [1].
Meta’s CacheLib uses hash index for SSD-resident objects by directly hashing each object to a 4KB SSD LBA (logical block address) block without an in-memory hash table. This makes CacheLib subject to a high SSD write amplification: each object insert/update invokes re-writing a 4KB LBA block. As a variant of CacheLib, Kangaroo [46] applies write-ahead log (WAL) to amortize the SSD write cost by buffering multiple cache objects hashed to the same LBA. Although it can reduce the SSD write amplification, the storage and management of WAL introduce non-negligible overhead in terms of SSD capacity and CPU/memory usage.
2.3 In-Storage Transparent Compression
@@
Figure 3: An illustration of (a) an SSD with built-in transparent compression, and (b) the expanded LBA space.
Figure 4: Overview of ZipCache architecture that employ three B+ trees to manage the DRAM cache tier, SSD cache tier, and large-size cache objects respectively.
3 Design
@@ 3.1 Architecture Overview
ZipCache is optimized for handling massive amount of small key-value items, which is not only practically important [18] but also poses significant challenges. ZipCache categorizes cache objects into three different size classes tiny, medium, and large by using pre-defined thresholds (e.g., 128B and 2KB). The three types of key-value items are handled differently: Tiny- and medium-size objects are stored across the DRAM and SSD tiers, while large-size objects are always SSD-resident, which is for maximizing the DRAM cache tier hit ratio. We compress in-memory tiny-size objects together in the unit of tree pages, and compress each in-memory medium-size object individually. In order to manage the different key-value items in DRAM and SSD cache tiers, ZipCache maintains three B+ trees:
  1. 1.
@@
3.2 DRAM Cache Tier
@@
To realize decompression early termination, we must be able to know which sub-page ${\bf p}_{i}$ contains the requested cache object before performing decompression, which however is impossible if we construct B+ tree leaf pages with conventional practice [33]. To address this issue, we construct each B+ tree leaf page in a hash-based manner as illustrated in Fig. 5: Let $\mathcal{K}$ denote the cache object key space and define a hash function $f:\ \mathcal{K}\to[1,\ n]$. For cache objects that fall into one B+ tree leaf page, we use the hash function $f$ to calculate their destined sub-pages inside the page. Therefore, to fetch a cache object from a B+ tree leaf page, we can determine its associated sub-page through a simple hashing and hence accordingly configure the decompression early termination.
Figure 5: Illustration of hash-based mapping between cache objects and sub-pages in a leaf page of $BT_{DRAM}$, including the leaf page decompression early termination.
@@
3.3 SSD Cache Tier
The emerging computational SSD provides built-in transparent compression [8], which brings multiple technical advantages. It not only offloads the resource-demanding (de)compression operations from the host CPUs, but also relieves the cache system from the complexities of handling the storage of variable-length post-compression data blocks.
@@ 3.3 SSD Cache Tier
Figure 6: Illustration of hash-based mapping between cache objects and 4KB sub-pages in a leaf page of $BT_{SSD}$, which decouples the the SSD read/write amplification from B+ tree leaf page size.
@@
Because both DRAM and SSD cache tiers use B+ tree index, we propose to apply page-oriented DRAM-to-SSD cache objects eviction to reduce the host-side write amplification $WA_{host}$. We choose the classical second-chance eviction policy for the purpose of implementation simplicity. Since B+ trees sort all the DRAM/SSD-resident cache objects based on their keys, the key range of a DRAM cache B+ tree leaf page may overlap with only one or few SSD cache B+ tree leaf pages. Unlike evicting small key-value items from random locations in cache, by evicting cold cache data in the unit of leaf pages, it only incurs the read-modify-write operations over a small number of SSD LBAs, leading to a small host-side write amplification $WA_{host}$. Moreover, thanks to the abundant spatial locality in real-world workloads, such page-based eviction is highly efficient, compared to object-based eviction as in CacheLib [18]. As demonstrated in prior work [46] and our experiments in Section 4, object-based eviction suffers from very high SSD write amplification.
@@
Figure 7: Comparison of throughput and overall DRAM/SSD cache hit ratio when the active working set is much larger than the DRAM/SSD cache capacity. We fixed the total active working set size $C_{WS}$ as 6TB and, under each category of workload locality, considered four different settings of {$C_{DRAM}$, $C_{SSD}$}: {64GB, 0.5TB}, {64GB, 1TB}, {128GB, 0.5TB}, and {128GB, 1TB}.
Figure 8: Comparison of throughput and DRAM cache hit ratio when the active working set fits into the DRAM/SSD cache (hence the SSD cache tier hit ratio is 100%). We fixed the DRAM cache tier capacity $C_{DRAM}$ as 64GB and considered $C_{DRAM}:C_{WS}$ ratio of 1:16, 1:8, 1:4, and 1:2.
3.4 Major Operations

### v2

@@ Abstract.
Key-Value Cache, Data Compression, DRAM/SSD Cache, Computational SSD
journalyear: 2024copyright: acmlicensedconference: The International Symposium on Memory Systems; September 30-October 3, 2024; Washington, VA, USAbooktitle: The International Symposium on Memory Systems (MEMSYS ’24), September 30-October 3, 2024, Washington, VA, USAdoi: 10.1145/3695794.3695805isbn: 979-8-4007-1091-9/24/09ccs: Information systems Database design and models
1. Introduction
@@
Key-value cache plays a crucial role in providing high-throughput, low-latency data services. Major Internet service providers, such as Google and Meta, often deploy a fleet of cache servers as the first line of defense to handle a massive influx of requests for key-value data, a type of unstructured data organized in simple forms as keys and values (e.g., “User ID” and “User name”). Key-value caching can accelerate data retrievals and alleviate the traffic to backend databases by serving from high-speed storage medium, typically DRAM, flash memory, or a combination of both, such as Meta’s CacheLib (Cac, 2024) and Ximalaya’s xcache (xca, 2022).
@@
The effectiveness of key-value caching hinges on its ability of accommodating the requested data in cache. While a larger cache intuitively leads to improved performance, expanding the capacity of key-value caches within data centers is often not simply a matter of hardware upgrade. Many practical constraints, such as server costs, cooling expenses, rack space, real estate limitations, and even human resource expenses must be taken into consideration. Let us consider the hardware cost as an example: Microsoft Azure reports that DRAM constitutes 50% of their server costs (Morgan, 2020), and Meta reports a similar trend (40% of the rack cost) (Maruf et al., 2023); Although the prices of high-speed NVMe SSDs are comparatively lower, they are still rather substantial (White, 2023; Olšan, 2023; Ao, 2023). Relying solely on hardware investment to increase cache capacity is apparently not a sustainable, cost-effective approach to keep up with the rapid growth of data. This poses an increasingly severe challenge in today’s data centers.
@@
A potential solution is compression. By condensing data to occupy a smaller footprint, one could virtually expand the cache capacity, allowing cache to hold more data, which in turn increases the cache hit ratio. Despite the adoption of data compression in computing systems in prior studies (Choukse et al., 2018; Ekman and Stenstrom, 2005; Pekhimenko et al., 2013; Zhao et al., 2015), interestingly, this simple idea has not gained much traction in key-value cache design. We believe that this lack of adoption in practice is due to several unique and critical issues inherent in key-value cache systems:
@@
$\bullet$ Issue #1: The commonly used hash indexing causes random, compression-unfriendly data placement. Most key-value cache systems adopt a hash index based structure to manage the key-value data (Red, 2023; Lim et al., 2014; Chandramouli et al., 2018). With hash indexing, keys are randomly dispersed in a flat, shallow data structure, which is advantageous for quick search in a large key space but comes with a detrimental effect for compression: Due to the nature of hash functions, the keys are evenly distributed, leaving unrelated data randomly mingled together. Such data layout is inherently difficult for effective data compression, as compression algorithms heavily rely on organizing similar data content within a close proximity.
@@
$\bullet$ Issue #2: The structure designed for managing small-size key-value items introduces a severe read/write amplification problem. Key-value workloads are known to be dominated by small-size data items. According to a study from Meta/Facebook, the majority of key-value items are (much) smaller than 500 bytes (Atikoglu et al., 2012). Since compressing each individual small key-value item yields limited or no benefits in size reduction, achieving effective compression requires to pack a collection of small key-value items for a reasonable compression ratio. However, this “optimization” would result in a substantial, undesirable increase in access operations, i.e., read/write amplification, when reading or updating a small key-value item in a much larger compression unit.
@@
We have implemented a prototype of ZipCache, which is a hybrid key-value cache with two cache layers, a DRAM layer and a flash memory layer. We use ScaleFlux’s CSD3000 SSD (Sca, 2024) with hardware acceleration for transparent on-device compression. Our experiments show promising results. In comparison to the state-of-the-art solutions, including CacheLib (Berg et al., 2020) and xcache (xca, 2022) and Kangaroo (McAllister et al., 2021), our evaluation results demonstrate that, with a design carefully optimized for data compression, ZipCache can achieve up to 72.4% higher throughput and 42.4% lower 90-percentile read latency, and reduces the SSD write amplification by up to 26.2$\times$. It is our hope that this work will motivate more future research to explore the full potential of the long-overlooked block compression for performance-critical caching systems.
@@
2. Background and Motivation
2.1. Data Compression
General-purpose block compression is realized by deduplicating repeated byte strings in a data block, referred to as LZ search (Ziv and Lempel, 1977, 1978). Although CPU-based LZ search suffers from low speed due to the high CPU cache miss rate, its reverse process (i.e., LZ decompression) can be much faster. The compression block size affects the trade-off between compression ratio [^1] In this work, we define compression ratio as $\frac{S_{orig}}{S_{comp}}\geq 1$, where $S_{orig}$ and $S_{comp}$ denote the size of the original and compressed data blocks. and (de)compression speed. Using the file samba in Silesia corpus (sil, 2018) as test data, Fig. 1 shows the LZ4 compression ratio and (de)compression latency under different block sizes. It shows about 4$\times$ speed performance difference between compression and decompression. As the compression block size continues to increase, the compression ratio first significantly improves and then gradually saturates.
Figure 1. Comparison of compression ratio, (de)compression latency under different compression block size.
The decompression process scans through the LZ-compressed byte stream to sequentially reconstruct the original data block. In theory, this process could terminate at any byte location, leading to a partially reconstructed data block. This makes it possible to realize early termination of decompression: Suppose LZ search compresses an $n$-segment data block ${\bf D}=[b_{1},b_{2},\cdots,b_{n}]$ into ${\bf C}$. When decompressing ${\bf C}$, the original data $[b_{1},\cdots,b_{k}]$ are successively reconstructed as $k$ grows from 1 to $n$. Let $\tau$ denote the latency of decompressing ${\bf C}$ to reconstruct the entire ${\bf D}$. If the decompression process terminates once after the first $m$ segments $[b_{1},\cdots,b_{m}]$, where $m\leq n$) have been reconstructed, we define its early termination factor $\gamma(m,n)=m/n$, and the decompression latency is about $\gamma(m,n)\cdot\tau$. Suppose we are only interested in obtaining one cache object in the $m$-th segment, we could reduce the latency by $1-m/n$ via decompression early termination. Using file samba in Silesia corpus (sil, 2018) as test data, we partitioned each 4KB data block into 256B ($n=16$) segments and measured the average LZ4 decompression latency under different early termination factor $\gamma(m,n)$ as shown in Table 1, which reveals substantial performance benefits.
Table 1. Decompression latency under different $\gamma(m,n)$.
@@ 2.1. Data Compression
2.2. Cache Index Data Structure
Most in-memory data stores (e.g., Redis (Red, 2023), FASTER (Chandramouli et al., 2018), MICA (Lim et al., 2014)) use hash index to reduce the latency and simplify the implementation. In contrast, most storage-based data stores (e.g., RocksDB (Roc, 2023; Dong et al., 2021), WiredTiger (Wir, 2024), Bw-tree (Levandoski et al., 2013)) employ tree index to reduce the index memory usage and embrace the storage block I/O interface. As for hybrid-DRAM/SSD caches, Meta’s CacheLib (Berg et al., 2020; Cac, 2024) uses hash index for the DRAM and SSD tiers, while Redis-compatible xcache (xca, 2022) (developed based on Redis (Red, 2023)) and Pika/RocksDB (pik, 2024) employs hash index for DRAM tier and log-structured merge (LSM) tree (O’Neil et al., 1996) index for SSD tier.
The index structure has a substantial effect on the efficacy of data compression. Since compression ratio is proportional to the byte content similarity, a tree index that sorts all the cache objects based on their keys is clearly more beneficial, in comparison to hash index that randomly hashes cache objects into data blocks. For the purpose of demonstration, Fig. 2(a) shows the 4KB block compression ratio under hash index and B+ tree index. The key is 8B unix timestamp and object size ranging from 4B to 64B, which are both extracted from the Bitstamp Exchange Data (BTC, 2023). By keeping objects with adjacent keys together, B+ tree achieves over 2$\times$ higher compression ratio than hash index. Although B+ tree index has a longer traversing latency than hash index, decompression tends to dominate the overall data access latency, which largely reduces the overall latency gap between B+ tree and hash index as shown in Fig. 2(b). A strong implication is that, although prior work on in-memory cache design widely adopted hash index, integrating block compression into in-memory cache makes tree index a favorable choice.
Figure 2. (a) Compression ratio of 4KB blocks under hash index and B+ tree index, and (b) latency of index traversing and block decompression under different total number of cache objects (hence different B+ tree depth), where key and values are obtained from Bitstamp Exchange Data (BTC, 2023).
Meta’s CacheLib uses hash index for SSD-resident objects by directly hashing each object to a 4KB SSD LBA (logical block address) block without an in-memory hash table. This makes CacheLib subject to a high SSD write amplification: each object insert/update invokes re-writing a 4KB LBA block. As a variant of CacheLib, Kangaroo (McAllister et al., 2021) applies write-ahead log (WAL) to amortize the SSD write cost by buffering multiple cache objects hashed to the same LBA. Although it can reduce the SSD write amplification, the storage and management of WAL introduce non-negligible overhead in terms of SSD capacity and CPU/memory usage.
2.3. In-Storage Transparent Compression
@@
Figure 3. An illustration of (a) an SSD with built-in transparent compression, and (b) the expanded LBA space.
Figure 4. Overview of ZipCache architecture that employ three B+ trees to manage the DRAM cache tier, SSD cache tier, and large-size cache objects respectively.
3. Design
@@ 3.1. Architecture Overview
ZipCache is optimized for handling massive amount of small key-value items, which is not only practically important (Berg et al., 2020) but also poses significant challenges. ZipCache categorizes cache objects into three different size classes tiny, medium, and large by using pre-defined thresholds (e.g., 128B and 2KB). The three types of key-value items are handled differently: Tiny- and medium-size objects are stored across the DRAM and SSD tiers, while large-size objects are always SSD-resident, which is for maximizing the DRAM cache tier hit ratio. We compress in-memory tiny-size objects together in the unit of tree pages, and compress each in-memory medium-size object individually. In order to manage the different key-value items in DRAM and SSD cache tiers, ZipCache maintains three B+ trees:
  1. (1)
@@
3.2. DRAM Cache Tier
@@
To realize decompression early termination, we must be able to know which sub-page ${\bf p}_{i}$ contains the requested cache object before performing decompression, which however is impossible if we construct B+ tree leaf pages with conventional practice (Graefe et al., 2011). To address this issue, we construct each B+ tree leaf page in a hash-based manner as illustrated in Fig. 5: Let $\mathcal{K}$ denote the cache object key space and define a hash function $f:\ \mathcal{K}\to[1,\ n]$. For cache objects that fall into one B+ tree leaf page, we use the hash function $f$ to calculate their destined sub-pages inside the page. Therefore, to fetch a cache object from a B+ tree leaf page, we can determine its associated sub-page through a simple hashing and hence accordingly configure the decompression early termination.
Figure 5. Illustration of hash-based mapping between cache objects and sub-pages in a leaf page of $BT_{DRAM}$, including the leaf page decompression early termination.
@@
3.3. SSD Cache Tier
The emerging computational SSD provides built-in transparent compression (Sca, 2024), which brings multiple technical advantages. It not only offloads the resource-demanding (de)compression operations from the host CPUs, but also relieves the cache system from the complexities of handling the storage of variable-length post-compression data blocks.
@@ 3.3. SSD Cache Tier
Figure 6. Illustration of hash-based mapping between cache objects and 4KB sub-pages in a leaf page of $BT_{SSD}$, which decouples the the SSD read/write amplification from B+ tree leaf page size.
@@
Because both DRAM and SSD cache tiers use B+ tree index, we propose to apply page-oriented DRAM-to-SSD cache objects eviction to reduce the host-side write amplification $WA_{host}$. We choose the classical second-chance eviction policy for the purpose of implementation simplicity. Since B+ trees sort all the DRAM/SSD-resident cache objects based on their keys, the key range of a DRAM cache B+ tree leaf page may overlap with only one or few SSD cache B+ tree leaf pages. Unlike evicting small key-value items from random locations in cache, by evicting cold cache data in the unit of leaf pages, it only incurs the read-modify-write operations over a small number of SSD LBAs, leading to a small host-side write amplification $WA_{host}$. Moreover, thanks to the abundant spatial locality in real-world workloads, such page-based eviction is highly efficient, compared to object-based eviction as in CacheLib (Berg et al., 2020). As demonstrated in prior work (McAllister et al., 2021) and our experiments in Section 4, object-based eviction suffers from very high SSD write amplification.
@@
3.4. Major Operations